500 people go to a concert, x paid €35 and the rest (y) paid €45. Total receipts were €20.500. Find x

Answer by checkley79(3341)

X+Y=500 MULTIPLY BY -35 & ADD.
35X+45Y=20,500
-35X-35Y=-17,500
------------------
10Y=3,000
Y=3,000/10
Y=300 ANS. FOR THE NUMBER OF PEOPLE WHO PAID 35.
X+300=500
X=500-300
X=200 ANS. FOR THE NUMBER OF PEOPLE WHO PAID 45.
PROOF:
35*200+45*300=20,500
7,000+13,500=20,500
20,500=20,500